The reaction 2 NO(g) = N2(g) + O2(g) has a Kc value of 2400 at 2000 K. If 0.850 M each of N2 and O2 are initially present in a 3.00-L vessel, calculate the equilibrium concentrations of No, N2 and O2.

1
Expert's answer
2022-06-02T08:46:02-0400

Kc = 2400

CM (N2) = CM (O2) = 0,85 mol/l


2NO = N2 + O2

Concentration of equilibrium:

NO = 2X

N2 = 0,85 - X

O2 = 0,85 - X


"K_c = \frac {[N_2][O_2]}{[NO]²}"


"2400= \frac {[0,85-x][0,85-x]}{[2x]²}"


"9600x\u00b2 = 0,7225 -1,7x +x\u00b2"


"9599x\u00b2 + 1,7x - 0,7225 = 0"


"X_{1,2} = \cfrac{-b ±\sqrt{b² -4ac} }{2a}"


"X_{1,2} = \frac{-1,7 ±\sqrt{1,7² -4*9599*(-0,7225)} }{2*9599}"


X= 0,00859

CM (N2) = CM (O2) = 0,85 - 0,00859 = 0,84141 mol/l

CM (NO) = 0,00859 * 2 = 0,01718 mol/l


ANSWER:

CM (N2) = 0,84141 mol/l

CM (O2) = 0,84141 mol/l

CM (NO) = 0,01718 mol/l